# Fingerprinting Implementations of Cryptographic Primitives and Protocols that Use Post-Quantum Algorithms

### v3

@@ 3.4 Digital Signatures
In this section, we present our findings for digital signatures following the first three research questions, as there are no hybrid digital signatures algorithms.
RQ1: Classical vs PQ classification. The datasets comprised 53976 samples, with 26976 from post-quantum digital signature schemes and 27000 from classical algorithms. Both Dataset1 and Dataset2 achieved 100% accuracy across all models. By reducing the feature set to the four memory-based features used in key exchange classification, we maintained similar accuracy levels. These results indicate that classical and PQ signature schemes are readily distinguishable based on their memory imprints.
@@
Table 11: Digital signatures, classical vs PQ, Dataset1 and Dataset2
RQ2: Multi-class PQ algorithm classification. The datasets comprised 48,000 samples, with 12,000 samples from each post-quantum digital signature scheme, ensuring equal representation across all variants. Ensemble learning models, particularly Random Forest and XGBoost, again led in performance, though with lower accuracy compared to key exchange classification. For Dataset2, Random Forest and XGBoost achieved 85% and 86% overall accuracy, respectively, which improved to 88% in Dataset1. These findings suggest that while classifying digital signature schemes may be more complex than key exchange algorithms using the same features, the results remain promising, particularly when computational costs are not influenced by additional load. Detailed metrics are provided in Tables 12 and 13.
@@
Table 12: Digital signatures, PQ Algorithms, Dataset1
@@
Table 13: Digital signatures, PQ Algorithms, Dataset2
RQ3: liboqs vs CIRCL, for Dilithium. The datasets consisted of 18,000 samples, with 9,000 samples each from liboqs and CIRCL implementations. Ensemble learning models, as anticipated, demonstrated superior performance. For Dataset1, we achieved 100% accuracy using only 4 features, and 95-96% accuracy for Dataset2, which increased to 100% when all features were included. Detailed results are presented in Tables 14 and 15. These findings indicate that the implementations are sufficiently distinct to reveal their corresponding libraries.
@@
Table 14: Digital signatures, liboqs vs CIRCL, Dataset1
@@
Table 15: Digital signatures, liboqs vs CIRCL, Dataset2
@@ 3.5 Discussion
Mitigation techniques. A potential defense against library fingerprinting involves restricting the visibility of memory usage data from processes spawned by one user to others, thereby mitigating unauthorized access to sensitive process metrics. Experimental observations have demonstrated that standard system monitoring tools can be exploited to locate a process’s ID and subsequently access its memory usage file, even when the process operates under elevated privileges. Concealing process identifiers may therefore serve as an effective countermeasure. In addition, memory obfuscation techniques, such as randomized memory allocation—which allocates buffers at random addresses or introduces dummy allocations to maintain a uniform usage profile—and the application of Oblivious RAM schemes to conceal access patterns, offer further protection, albeit with notable performance overheads and primarily as a subject of ongoing research.
@@
4 Fingerprinting of Protocols
In this section we describe our results on fingerprinting key exchange algorithms used by a protocol. We seek to answer the following questions:
@@ 4.1 Methodology
Device Specifications All experiments were conducted across four different devices. The first device was configured with Ubuntu 22.04, running on an Intel Core i5-11400H CPU, featuring 6 physical cores and 12 virtual cores, alongside 16 GB of RAM. The second and third devices were both operating on Windows; the second device was powered by an Intel Core i5-6300U CPU, with 2 physical cores and 4 virtual cores, and the third device utilized an Intel Core i7-8565U CPU, offering 4 physical cores and 8 virtual cores, both equipped with 16 GB of RAM. The final device was running macOS Sonoma, powered by the Apple M2 chip, with 16 GB of RAM.
Datasets For local traffic generation and data collection, we utilize packet sniffing tools such as Wireshark and tcpdump. Additionally, we supplement our dataset with publicly available packet captures from sources such as Cloudflare[73] and Wireshark[1]. For local data collection, to emulate real world conditions, we simultaneously run the standard and post quantum implementations so that our packet captures contain not only connections that use classical algorithms but also post quantum.
4.2 Fingerprinting Method
We aim to identify the key exchange algorithm used for a successful connection between two entities over a network where a successful connection means that the same key exchange algorithm was used on both entities. To ensure this, the desirable scenario is to have packets from both entities but if we do not, we also provide a solution to the issue.
We rely just on the clearly transmitted key sizes embedded in the packets to identify and classify the key exchange algorithm used for connection initiation. Passing the packet capture data through our protocol specific filter provides us with a list of key exchange algorithms used for connections within that protocol.
Fingerprinting TLS, OpenVPN, QUIC, OIDC
A single approach suffices to identify the key exchange algorithm used for these protocols because they all use some implementation of TLS to ensure secure key exchange. To identify the key used by a connection we need to extract it from the tls layer of a packet which vary across TLS 1.2 and TLS 1.3.
@@ 4.3 Results
RQ1: Classical vs PQ protocols. For SSH connections, on Ubuntu, MacOS and Windows we successfully identify all classical and post quantum connections. Since OpenSSH(version>=9.0) supports only 1 hybrid algorithm by default, all post quantum connections used that key exchange. As for TLS connections we achieved complete success for connections between devices running Ubuntu and MacOS. Our fingerprinting approach is able to detect and identify key exchanges for most QUIC connections. All TLS connections for key exchange within entities of OIDC such as the relying party, user agent and OIDC provider were also successfully identified by our approach. Based on our findings we can say that classical and post quantum algorithms are easily distinguishable from their key sizes as the post-quantum public keys and cipher texts tend to be exponentially larger than their classical counterparts.
@@ 4.4 Discussion
Attacker capability to conduct the fingerprinting. An adversary on the same network as the victim can intercept handshake packets exchanged between the victim and external servers, allowing them to determine whether the connection employs post-quantum key exchange. Since key sizes and public keys are transmitted in plaintext within the ClientHello and ServerHello messages, the adversary can analyze these parameters to infer the cryptographic scheme in use. This exposure enables the identification of connections relying on vulnerable key exchange algorithms, potentially allowing the adversary to exploit weaknesses and compromise security.
Defenses
A potential mitigation for protocol fingerprinting is to ensure that key shares used in deriving a shared secret are not transmitted in plaintext over the network. Although the Encrypted Client Hello (ECH) effectively encrypts the client’s key share—thereby preventing its exposure—this approach does not extend to the server, whose key share remains transmitted in clear during the Server Hello, thereby presenting a vulnerability that could facilitate fingerprinting attacks.
Limitations
Our approach misses some TLS connections from Windows machines because of custom headers. Although our approach is also able to handle fragmented packets in most cases, fragmentation of client and/or server hello packets in TLS 1.2 for OpenVPN connections fails to be recognized by it.
5 Fingerprinting SNARK libraries
In this section, we describe the fingerprinting of SNARK generation libraries. We seek to answer the following question:
@@ 5.1 Methodology
Libraries we consider. We investigate two SNARK libraries: lattice-zkSNARK [38], which integrates post-quantum cryptographic techniques within the libsnark framework [2], and pysnark [3], which relies on the classical libsnark backend. For a rigorous comparative analysis, both libraries were evaluated by generating SNARKs for a standardized program.
@@
Datasets. CPU cycle count and memory usage during SNARK generation and verification were used as classification features. CPU cycle counts were obtained using the perf command, and memory usage was measured from /proc/[pid]/status over 1000 experimental runs. Two datasets were generated: Dataset1, which minimized background processes, and Dataset2, which incorporated CPU-intensive tasks such as matrix multiplications, prime number calculations, large list sorting, and hash computations. While these additional tasks influenced computation time, they did not affect memory usage for the evaluated processes. Owing to limited data availability, the final dataset used in this study comprised a combination of Dataset1 and Dataset2.
5.2 Fingerprinting Method
Our initial analysis reveals significant differences between the two libraries in terms of computational cycles and memory usage, enabling clear library-specific fingerprinting. For example, as seen from Fig 1 and 2, the PQ lattice-zkSNARK have substantially higher execution times compared to the classical pysnark, reflecting the additional cryptographic complexity inherent in post-quantum systems. Moreover, the distinct cycle distributions across CPU cores and unique memory footprints associated with each library further reinforce their identifiable characteristics. These findings underline the potential for accurate differentiation of SNARK implementations, offering valuable insights for security analysis and optimization in both post-quantum and classical SNARK contexts.
Based on these results, we formulate the fingerprinting problem as a machine learning binary classification, with classes identified as classical and post-quantum. We start the classification task with 19 features, of which 12 features represent the cycles used by each core of the CPU and remaining 7 features represent a process’s memory usage, including total virtual memory, physical memory in use, data segment, stack, executable code, shared libraries, and page table entries. We narrow down our feature set to 2 CPU cycle based features using the Chi-square statistical test as the scoring function to get the top features by importance. Finally, we use an 80-20 train-test split of the dataset.
@@ 5.3 Results
The dataset comprised of 4000 samples, evenly split between pysnark  and lattice_zksnark . The XGBoost model achieved 100% classification accuracy. The results are presented in Table 16. This level of accuracy can be extrapolated to larger datasets, given the distinctive and consistent patterns observed in the computational and memory usage metrics.
@@
Table 16: SNARK libraries, pysnark vs lattice_zksnark, Dataset1 & Dataset2
@@ 6 Case Studies
In this section we describe how fingerprinting post-quantum libraries and protocols can benefit real world applications.
@@ 6.1 Quartz Integration
As part of the real-world transition to PQC [53] enterprises must be able to assess their PQC-readiness. To support these efforts, Cisco has released as open source a research prototype called "Quartz" - Quantum Risk and Threat Analyzer [19]. Quartz supports static and dynamic scanning of several protocols and processes, analyzes and identifies quantum-vulnerable cryptography protocols and libraries being used, and carries out risk analysis that can help in remediation of such vulnerabilities using quantum-resistant algorithms and implementations such as PQC standards. It scans and analyzes network communications, cloud applications, databases, operating systems, file systems. Quartz also analyzes source code, SQL queries, and cloud account activities.
Quartz relies on communication meta-data and payload if it can access, the configurations and code/scripts in order to determine the type of cryptography protocols, but it does not have the ability to analyze performance metrics and fingerprinting capabilities as proposed in this paper. Our fingerprinting techniques can be used by a system like Quartz in order to be able to identify quantum-vulnerable cryptography protocols and libraries across the communication and computing stack in a black-box and/or white-box manner.
Integration with Quartz. We have enhanced the Quartz system by incorporating advanced features for post-quantum connection detection in TLS and the classification of cryptographic key exchange and digital signature algorithms based on CPU cycle and memory usage. Furthermore, we have developed and integrated dedicated API endpoints into the Quartz system, facilitating seamless backend execution of the relevant scripts. We will make our changes available in the main Quartz repository to promote transparency and accessibility.
@@ 6.2 Identifying Post Quantum TLS Connections on domains from Tranco
To identify servers on the internet using PQC keys for TLS connections, we conducted a targeted probe of 1 million domains selected from the Tranco dataset [43]. This effort yielded 160,023 connection attempts to these domains, from which 4,988 unique IP addresses were identified as potentially supporting PQ keys. It is worth noting that not all domains could be probed due to DNS resolution failures, impacting the total connection attempts.
Occasionally, our probing process generated multiple algorithmic suggestions per connection, including both classical and PQ methods. For instance, the IP associated with the domain 1-800-FLOWERS.com suggested the use of either ECDH-P384 (classical) or Classic-McEliece-348864 (PQ) key, but manual validation confirmed the use of ECDH-P384. While these minor discrepancies introduce a small margin of error, we maintain that a slight overestimation of PQC presence is preferable to underdetection in this exploratory context. Table 17 presents the top 10 organizations ranked by the number of IP addresses associated with organisations.
@@
Table 17: Top 10 Organizations by Number of IPs
@@ 7 Related Work
Side-channel attacks. The prevalence of side-channel attacks on classical cryptographic algorithms, such as RSA, AES, and ECC, provides a strong precedent for the potential occurrence of similar attacks on PQ algorithms. This work [76] revealed that side-channel attacks like simple power analysis can potentially reveal the secret key through power consumption variations during the execution of ECC point operations. The study [42] showed that Differential Power Analysis can be used to extract AES keys by analyzing power consumption patterns during encryption operations. This paper [29] presents an acoustic side-channel attack that can extract RSA keys by analyzing the sound produced by the computer during cryptographic operations. The study in [16] showed that they can extract private keys from an OpenSSL-based web server running on a machine in the local network by timing side-channel attack against OpenSSL. CRYSTALS-Kyber implementation has been broken by  [25] (after it was selected as a finalist and before the standardization).
Traffic fingeprinting. In this work [78] the authors show a novel approach to protocol identification in encrypted traffic by utilizing observable features such as packet timing, size, and direction, achieving high accuracy without relying on packet contents or host information. The study [77] proposes an end-to-end encrypted traffic classification method with only a one-dimensional convolution neural network.
Benchmarking PQ protocols. The work [67] introduces a framework for benchmarking PQ cryptographic algorithms within TLS, specifically analyzing the impact of latency and packet loss on hybrid key exchanges and digital signatures.

### v4

@@ 3.4 Digital Signatures
In this section, we present our findings for digital signatures following the first three research questions, as there are no hybrid digital signature algorithms.
RQ1: Classical vs PQ classification. The datasets comprised 53976 samples, with 26976 from post-quantum digital signature schemes and 27000 from classical algorithms. Both Dataset1 and Dataset2 achieved 100% accuracy across all models. By reducing the feature set to the four memory-based features used in key exchange classification, we maintained similar accuracy levels. These results indicate that classical and PQ signature schemes are readily distinguishable based on their memory imprints. Detailed metrics are provided in Table XI.
@@
TABLE XI: Digital signatures, classical vs PQ, both datasets
RQ2: Multi-class PQ algorithm classification. The datasets comprised 48,000 samples, with 12,000 samples from each post-quantum digital signature scheme, ensuring equal representation across all variants. Ensemble learning models, particularly Random Forest and XGBoost, again led in performance, though with lower accuracy compared to key exchange classification. For Dataset2, Random Forest and XGBoost achieved 85% and 86% overall accuracy, respectively, which improved to 88% in Dataset1. These findings suggest that while classifying digital signature schemes may be more complex than key exchange algorithms using the same features, the results remain promising, particularly when computational costs are not influenced by additional load. Detailed metrics are provided in Tables XII and XIII.
@@
TABLE XII: Digital signatures, PQ Algorithms, Dataset1
@@
TABLE XIII: Digital signatures, PQ Algorithms, Dataset2
RQ3: liboqs vs CIRCL, for Dilithium. The datasets consisted of 18,000 samples, with 9,000 samples each from liboqs and CIRCL implementations. Ensemble learning models, as anticipated, demonstrated superior performance. For Dataset1, we achieved 100% accuracy using only 4 features, and 95-96% accuracy for Dataset2, which increased to 100% when all features were included. Detailed results are presented in Tables XIV and XV. These findings indicate that the implementations are sufficiently distinct to reveal their corresponding libraries.
@@
TABLE XIV: Digital signatures, liboqs vs CIRCL, Dataset1
@@
TABLE XV: Digital signatures, liboqs vs CIRCL, Dataset2
@@ 3.5 Discussion
Mitigation techniques. A potential defense against library classification involves restricting the visibility of memory usage data from processes spawned by one user to others, thereby mitigating unauthorized access to sensitive process metrics. Experimental observations have demonstrated that standard system monitoring tools can be exploited to locate a process’s ID and subsequently access its memory usage file, even when the process operates under elevated privileges. Concealing process identifiers may therefore serve as an effective countermeasure. In addition, memory obfuscation techniques, such as randomized memory allocation—which allocates buffers at random addresses or introduces dummy allocations to maintain a uniform usage profile—and the application of Oblivious RAM schemes to conceal access patterns, offer further protection, albeit with notable performance overheads and primarily as a subject of ongoing research.
@@
4 Classification of Protocols
In this section we describe our results on the classification of key exchange algorithms used by a protocol. We seek to answer the following questions:
@@ 4.1 Methodology
Device specifications. All experiments were conducted across four different devices. The first device was configured with Ubuntu 22.04, running on an Intel Core i5-11400H CPU, featuring 6 physical cores and 12 virtual cores, alongside 16 GB of RAM. The second and third devices were both operating on Windows; the second device was powered by an Intel Core i5-6300U CPU, with 2 physical cores and 4 virtual cores, and the third device utilized an Intel Core i7-8565U CPU, offering 4 physical cores and 8 virtual cores, both equipped with 16 GB of RAM. The final device was running macOS Sonoma, powered by the Apple M2 chip, with 16 GB of RAM.
Datasets. For local traffic generation and data collection, we utilize packet sniffing tools such as Wireshark and tcpdump. Additionally, we supplement our dataset with publicly available packet captures from sources such as Cloudflare[qacafe_sample_captures] and Wireshark[wireshark_sample_captures]. For local data collection, to emulate real world conditions, we simultaneously run the standard and post quantum implementations so that our packet captures contain not only connections that use classical algorithms but also post quantum.
4.2 Classification Method
We aim to identify the key exchange algorithm used for a successful connection between two entities over a network where a successful connection means that the same key exchange algorithm was used on both entities. The desirable scenario is to have packets from both entities but if we do not, we also provide a solution to the issue.
We rely just on the clearly transmitted key sizes embedded in the packets to identify and classify the key exchange algorithm used for connection initiation. To achieve this, the packet capture data is passed through a protocol-specific filter. This filter is tailored to the structure of each protocol (e.g., TLS or SSH) and is responsible for selectively extracting only the key exchange–related messages, identified by their specific message types, while excluding all other protocol traffic. The result is a concise list of key exchange algorithms observed within each protocol.
Classification in TLS, OpenVPN, QUIC, OIDC
A single approach suffices to identify the key exchange algorithm used by these protocols, as they all rely on some implementation of the TLS protocol to establish secure key exchange. To determine the key exchange method used by a given connection, it is necessary to extract the relevant information from the TLS layer of the packet, which differs between TLS 1.2 and TLS 1.3. Notably, the handshake messages containing the key exchange information are transmitted in the clear—that is, they are not encrypted and are fully observable to any party monitoring the communication.
@@ 4.3 Results
[…]
@@ 4.4 Discussion
Attacker capability to conduct the classification. An adversary on the same network as the victim can intercept handshake packets exchanged between the victim and external servers, allowing them to determine whether the connection employs post-quantum key exchange. Since key sizes and public keys are transmitted in plaintext within the ClientHello and ServerHello messages, the adversary can analyze these parameters to infer the cryptographic scheme in use. This exposure enables the identification of connections relying on vulnerable key exchange algorithms, potentially allowing the adversary to exploit weaknesses and compromise security.
Defenses. A potential mitigation for protocol identification is to ensure that key shares used in deriving a shared secret are not transmitted in plaintext over the network. Although the Encrypted Client Hello (ECH) effectively encrypts the client’s key share—thereby preventing its exposure—this approach does not extend to the server, whose key share remains transmitted in clear during the Server Hello, thereby presenting a vulnerability that could facilitate identification.
Limitations. Our approach misses some TLS connections from Windows machines because of custom headers. Although our approach is also able to handle fragmented packets in most cases, fragmentation of client and/or server hello packets in TLS 1.2 for OpenVPN connections fails to be recognized by it.
4.5 Integration of PQ digital signatures in secure protocols
5 Classifying SNARK libraries
In this section, we describe the classification of SNARK generation libraries. We seek to answer the following question:
@@ 5.1 Methodology
Libraries we consider. We investigate two SNARK libraries: lattice-zkSNARK [ISW21], which integrates post-quantum cryptographic techniques within the libsnark framework [libsnark], and pysnark [pysnark], which relies on the classical libsnark backend. For a rigorous comparative analysis, both libraries were evaluated by generating SNARKs for a standardized program.
@@
Datasets. CPU cycle count and memory usage during SNARK generation and verification were used as classification features. CPU cycle counts were obtained using the perf command, and memory usage was measured from /proc/[pid]/status over 1000 experimental runs. Two datasets were generated: Dataset1_SNARK , which minimized background processes, and Dataset2_SNARK , which incorporated CPU-intensive tasks such as matrix multiplications, prime number calculations, large list sorting, and hash computations. While these additional tasks influenced computation time, they did not affect memory usage for the evaluated processes. Owing to limited data availability, the final dataset used in this study comprised a combination of Dataset1_SNARK and Dataset2_SNARK .
5.2 Classification Method
Our initial analysis reveals significant differences between the two libraries in terms of computational cycles and memory usage, enabling clear library-specific classification. For example, as seen from Fig 3 and 4, the PQ lattice-zkSNARK have substantially higher execution times compared to the classical pysnark, reflecting the additional cryptographic complexity inherent in post-quantum systems. Moreover, the distinct cycle distributions across CPU cores and unique memory footprints associated with each library further reinforce their identifiable characteristics. These findings underline the potential for accurate differentiation of SNARK implementations, offering valuable insights for security analysis and optimization in both post-quantum and classical SNARK contexts.
Based on these results, we formulate the problem as a machine learning binary classification, with classes identified as classical and post-quantum. We start the classification task with 19 features, of which 12 features represent the cycles used by each core of the CPU and remaining 7 features represent a process’s memory usage, including total virtual memory, physical memory in use, data segment, stack, executable code, shared libraries, and page table entries. We narrow down our feature set to 2 CPU cycle based features using the Chi-square statistical test as the scoring function to get the top features by importance. Finally, we use an 80-20 train-test split of the dataset.
@@ 5.3 Results
The dataset comprised of 4000 samples, evenly split between pysnark  and lattice_zksnark . The XGBoost model achieved 100% classification accuracy. The results are presented in Table XVI. This level of accuracy can be extrapolated to larger datasets, given the distinctive and consistent patterns observed in the computational and memory usage metrics.
@@
TABLE XVI: SNARK libraries, pysnark vs lattice_zksnark, Dataset1_SNARK Dataset2_SNARK
@@ 6 Case Studies
In this section we describe how classification of post-quantum libraries and protocols can benefit real world applications.
@@ 6.1 Quartz Integration
As part of the real-world transition to PQC [nist-pqc-details] enterprises must be able to assess their PQC-readiness. To support these efforts, Cisco has released as open source a research prototype called ”Quartz” - Quantum Risk and Threat Analyzer [quartz]. Quartz supports static and dynamic scanning of several protocols and processes, analyzes and identifies quantum-vulnerable cryptography protocols and libraries being used, and carries out risk analysis that can help in remediation of such vulnerabilities using quantum-resistant algorithms and implementations such as PQC standards. It scans and analyzes network communications, cloud applications, databases, operating systems, file systems. Quartz also analyzes source code, SQL queries, and cloud account activities.
Quartz relies on communication meta-data and payload if it can access, the configurations and code/scripts in order to determine the type of cryptography protocols, but it lacks the ability to analyze performance metrics and classification capabilities as proposed in this paper. Our classification techniques can be used by a system like Quartz in order to be able to identify quantum-vulnerable cryptography protocols and libraries across the communication and computing stack in a black-box and/or white-box manner.
Integration with Quartz. We enhanced the Quartz system by incorporating advanced features for post-quantum connection detection in TLS and the classification of cryptographic key exchange and digital signature algorithms based on CPU cycle and memory usage. Furthermore, we have developed and integrated dedicated API endpoints into the Quartz system, facilitating seamless backend execution of the relevant scripts. We will make our changes available in the main Quartz repository to promote transparency and accessibility. We provide an anonymous repository with a cloned version of Quartz and our integrated features, available at https://anonymous.4open.science/r/Fingerprint-809D/.
@@ 6.2 Identifying Post Quantum TLS Connections on domains from Tranco
To identify servers on the internet using PQC keys for TLS connections, we conducted a targeted probe of 1 million domains selected from the Tranco dataset [LePochat2019tranco]. This effort yielded 160,023 connection attempts to these domains, from which 4,988 unique IP addresses were identified as potentially supporting PQ keys. It is worth noting that not all domains could be probed due to DNS resolution failures, impacting the total connection attempts.
Occasionally, our probing process generated multiple algorithmic suggestions per connection, including both classical and PQ methods. For instance, the IP associated with the domain 1-800-FLOWERS.com suggested the use of either ECDH-P384 (classical) or Classic-McEliece-348864 (PQ) key, but manual validation confirmed the use of ECDH-P384. While these minor discrepancies introduce a small margin of error, we maintain that a slight overestimation of PQC presence is preferable to underdetection in this exploratory context. Table XVII presents the top 10 organizations ranked by the number of IP addresses associated with organisations.
@@
TABLE XVII: Top 10 Organizations by Number of IPs
@@ 7 Related Work
Side-channel attacks. The prevalence of side-channel attacks on classical cryptographic algorithms, such as RSA, AES, and ECC, provides a strong precedent for the potential occurrence of similar attacks on PQ algorithms. Walter et al. [walter2004] revealed that side-channel attacks like simple power analysis can potentially reveal the secret key through power consumption variations during the execution of ECC point operations. Kocher et al. [kocher1999] showed that Differential Power Analysis can be used to extract AES keys by analyzing power consumption patterns during encryption operations. Genkin et al. [rsaAcoustic] presents an acoustic side-channel attack that can extract RSA keys by analyzing the sound produced by the computer during cryptographic operations. Brumley et al. [brumley2005remote] showed that they can extract private keys from an OpenSSL-based web server running on a machine in the local network by timing side-channel attacks against OpenSSL. Wang et al. [wang2023dvfs] further demonstrates that dynamic voltage and frequency scaling (DVFS) can cause data-dependent power consumption to influence CPU frequency, creating timing variations even in “constant-time” implementations and enabling remote key recovery and cross-component leakage. Crystals-Kyber implementations have also been broken by Dubrova et al. [dubrova2023breaking] (after it was selected as a finalist and before the standardization).
Traffic identification. Wright et al. [wright2006inferring] shows a novel approach to protocol identification in encrypted traffic by utilizing observable features such as packet timing, size, and direction, achieving high accuracy without relying on packet contents or host information. Wang et al. [1dcnn] proposes an end-to-end encrypted traffic classification method with only a one-dimensional convolution neural network. Xue et al. [xue2022] demonstrated how the network flow of OpenVPN could be fingerprinted by identifying fixed patterns within OpenVPN traffic. Similarly, Panchenko et al. [torf] presented an attack on Tor, revealing that traffic analysis could be used to identify visited websites, despite the use of encryption, by exploiting unique traffic patterns generated by different websites.
Benchmarking PQ protocols. Paquin et al. [PQCrypto:PaqSteTam20] introduces a framework for benchmarking PQ cryptographic algorithms within TLS, specifically analyzing the impact of latency and packet loss on hybrid key exchanges and digital signatures.